OpenStudy (anonymous):

How do i solve this please: 2 square root of x minus square root of 2x+1 = 1

OpenStudy (anonymous):

\[2\sqrt{x} - \sqrt{2x+1} = 1\]

OpenStudy (shamim):

take square of both side first

OpenStudy (anonymous):

i got 4x = 2x+1 +1

OpenStudy (anonymous):

and after solving that I got x=1

OpenStudy (shamim):

no

OpenStudy (shamim):

ok. can i give u an attachment

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you so much !

OpenStudy (shamim):

wc

OpenStudy (mathstudent55):

@shamin Great solution, but you forgot 1 step. When you square both sides of an equation, you must check all solutions in the original equation because the squaring of both sides may introduce extraneous solutions.  The solution is only x = 4